### v1

# Vision Language Model-based Testing of Industrial Autonomous Mobile Robots

Abstract
Autonomous Mobile Robots (AMRs) are deployed in diverse environments (e.g., warehouses, retail spaces, and offices), where they work alongside humans. Given that human behavior can be unpredictable and that AMRs may not have been trained to handle all possible unknown and uncertain behaviors, it is important to test AMRs under a wide range of human interactions to ensure their safe behavior. Moreover, testing in real environments with actual AMRs and humans is often costly, impractical, and potentially hazardous (e.g., it could result in human injury). To this end, we propose a Vision Language Model (VLM)-based testing approach (RVSG) for industrial AMRs developed by PAL Robotics in Spain. Based on the functional and safety requirements, RVSG uses the VLM to generate diverse human behaviors that violate these requirements. We evaluated RVSG with several requirements and navigation routes in a simulator using the latest AMR from PAL Robotics. Our results show that, compared with the baseline, RVSG can effectively generate requirement-violating scenarios. Moreover, RVSG-generated scenarios increase variability in robot behavior, thereby helping reveal their uncertain behaviors.
Keywords Autonomous robots  $\cdot$ Vision language models  $\cdot$ Human behaviors  $\cdot$ Simulation-based testing
1 Introduction
PAL Robotics (Spain) [1] is a leading service robot manufacturer, with deployments in diverse industries, including healthcare, retail, and agri-food. Notably, their Autonomous Mobile Robots (AMRs) excel at safe and efficient navigation in environments, e.g., offices and warehouses, automating tasks alongside humans. Unpredictable human behavior and frequent interactions between humans and robots place high demands on AMR navigation [2]. To ensure dependable operation, PAL Robotics defines key functional and safety (F&S) requirements for various deployments that AMRs must meet.
Scenario-based testing offers a systematic approach to test AMRs’ navigation under various human-robot interactions [3]. Integrating F&S requirements into automated testing helps generate realistic testing scenarios that violate these requirements [3]. Besides, to improve testing efficiency, a small set of test scenarios that capture diverse human behaviors shall be used, avoiding similar behaviors caused by the same faults [4]. Moreover, given the high cost of testing in real environments with actual robots and humans, the impracticality (e.g., asking humans to exhibit a specific behavior), and the potential safety risks, such as collisions that could damage robots or injure humans, simulation-based testing provides a safe and cheap alternative [5]. However, ensuring realistic human behaviors in simulation remains a significant challenge [5].
To address the challenge mentioned above, we use Vision Language Models (VLMs) to generate test scenarios, which may achieve human-level performance in commonsense reasoning about various tasks [6]. VLMs can extract key information from both images and text, enabling contextual understanding and drawing logical conclusions. Thus, we conjecture that VLMs are naturally well-suited to analyze textual requirements and exhibit strong capabilities in scenario interpretation and construction. This makes VLMs highly capable of analyzing F&S requirements and generating natural language descriptions of test scenarios and human behaviors. An often overlooked but crucial aspect of scenario realism is that human behavior varies depending on the environment. For example, typical human behaviors in warehouses include lifting, sorting packages, and interacting with shelves. In simulation-based testing, it is thus essential that modeled human behaviors remain consistent with the logical context of the given environment, which directly influences the configuration of such behaviors. VLMs, using their commonsense reasoning capability, can generate contextually appropriate human behaviors tailored to specific environments, which is difficult to achieve with traditional methods, e.g., search algorithms.
To this end, we propose RVSG, a requirement-driven test scenario generation method for robot navigation tasks based on VLMs having two stages: (i) environment preprocessing and (ii) test scenario generation. In the first stage, RVSG captures and labels images of the simulated environment, then uses the VLM to extract visual features to support scenario generation. In the second stage, the prompt generator adapts pre-designed natural language prompt templates based on the given requirement and task. Through multi-turn conversations, the VLM generates detailed scenario descriptions and human behavior configurations, which are decoded into executable human agents and run together with the navigation task in the simulator to obtain test results. These results are then fed back to the prompt generator to optimize the human configurations, improving test scenarios. Moreover, RVSG uses generated scenarios, human configurations, and feedback to guide the VLM in producing more diverse scenarios. We evaluated RVSG and the baseline method on the latest AMR from PAL Robotics using the Gazebo simulator with different requirements and navigation routes. We also employed multiple metrics to assess the performance and diversity of the generated test scenarios. Our results show that RVSG can generate more effective requirement-violating test scenarios with a certain degree of diversity, increasing the variability of robot behaviors that help reveal behavioral uncertainties. In addition, different navigation routes affect both the performance of the generated scenarios and the stability of scenario execution.
In summary, our contributions are: (1) a novel requirement-driven VLM-based test scenario generation method for AMRs; (2) a set of prompt templates that enable RVSG to understand the environment, generate test scenarios and human configurations, and dynamically incorporate simulation feedback and historical scenario information for improving scenario generation; (3) evaluate the test scenarios generated by RVSG with multiple metrics on the latest AMR from PAL Robotics within the Gazebo simulator, covering a variety of requirements and navigation routes.
2 Industrial Context
This work is conducted in the context of the RoboSAPIENS European project [7], which aims to build and test self-adaptive robots capable of operating safely in uncertain and unknown environments, with adaptation mechanisms implemented as software within the robots. PAL Robotics (Spain) [1] is one of the industrial use-case providers in RoboSAPIENS. The company is recognized for developing and deploying service robots across various domains, including retail, office, and warehouse environments. Their portfolio includes legged robots, social robots, inventory robots, and Autonomous Mobile Robots (AMRs) for intralogistics applications.
This study focuses on AMRs, which are deployed in environments such as offices and warehouses, where they navigate within a defined space and operate alongside humans. Naturally, the interaction between humans and AMRs must be safe. Typically, AMRs use two types of navigation algorithms: a global planner, which determines the optimal path to reach the destination, and a local planner, which accounts for dynamic obstacles in the environment (including humans) and adapts the path accordingly. To this end, an AMR is typically equipped with a set of LiDAR sensors to support various functions such as path planning, collision avoidance, and autonomous navigation. Fig. 1 shows the TIAGo OMNI Base AMR, which is used in many practical applications, both individually and as part of a fleet, to perform its intended operations collaboratively.
Within the RoboSAPIENS, self-adaptive behaviors are implemented using the MAPLE-K loop [8], an extension of the well-established MAPE-K loop for self-adaptation in software systems. As presented in Fig. 1, in the MAPLE-K loop, an AMR continuously monitors (Monitor) its environment using dedicated sensors, analyzes the data (Analyze), and, upon detecting a situation that requires adaptation, plans (Plan), validates (Legitimate), and executes (Execute) the selected adaptation. Throughout this process, all information becomes part of the robot’s knowledge base (Knowledge), which can be accessed at any stage of the loop. The Legitimate phase is newly introduced in MAPLE-K.
Given the diverse deployment environments of AMRs, it is essential to ensure that they navigate efficiently, stably, and safely while posing no threat to humans [5]. Functional and safety requirements are typically provided by domain experts, based on their domain knowledge and the specifications of both the AMRs and their operating environments. If a potential threat is detected, AMRs adapt their behavior, such as changing their path, stopping, or even communicating with nearby humans. However, due to unpredictable human behavior, it is unknown in advance how humans will interact in the environment where an AMR is deployed. Therefore, it is paramount to test AMRs under a variety of human behaviors, ensuring (to a certain degree) the dependability of their responses. Moreover, conducting tests in real-world environments with real humans is expensive and often impractical [5]. Instead, simulation-based testing offers an affordable alternative, which is the focus of this paper. In general, there is a clear need for systematic methods to test AMR behavior in simulation settings, where diverse human behaviors can be simulated to assess the dependability of AMRs. As Fig. 1 shows, within the context of the MAPLE-K loop, our simulation-based testing approach is designed to support three key phases of MAPLE-K: (1) it enables testing of the MAPLE-K loop implemented in AMRs by generating scenarios that force the AMR to trigger adaptations, thereby ensuring whether the Plan component performs its intended functionality; (2) it supports the creation of critical scenarios for the Legitimate component, helping assess whether selected adaptations are safe to execute; and (3) it contributes to building a knowledge base of critical scenarios for the Knowledge component, contributing to the iterative development of MAPLE-K versions.
Figure 1: A representative autonomous mobile robot from PAL Robotics, TIAGo OMNI Base. It is planned to be equipped with the MAPLE-K loop to enable self-adaptation in uncertain and unknown environments. The phases of MAPLE-K highlighted in orange are those supported by RVSG.
3 Method
3.1 RVSG Overview
Fig. 2 presents an overview of RVSG for generating scenarios during the iterative AMR testing. As shown in Fig. 2, RVSG has two main stages: (i) environment preprocessing and (ii) test scenario generation. In the first stage, RVSG achieves the Map Image Preprocessing and Environment Description Generation. At Map Image Preprocessing step ( 1 in Fig. 2), RVSG first obtains a top-down view of the simulation world as a map image. This map can then be labeled to highlight the spatial structure, facilitating subsequent layout analysis and task planning. Based on the labeled map, RVSG constructs a grid map with precise waypoint mapping to refine the layout and support robot navigation description and human configuration generation. Both processed images are provided to the VLM for Environment Description Generation at 2. Within 2, the VLM utilizes multi-turn conversations to extract key layout and navigation information from the images, and generates a natural language description of the environment, which assists in subsequent test scenario generation.
Figure 2: Overview of RVSG during the iterative testing of autonomous mobile robots.
For test scenario generation, at 3, RVSG integrates the specific F&S requirement, robot navigation route, and the preprocessed map images and environment description. These inputs are structured into carefully designed prompts (detailed in Fig. 3) for the VLM, guiding its automated understanding and analysis to generate human configurations. At 4, within the Human Configuration Generation by Multi-turn Conversation, Prompt Generator automatically selects appropriate initial multi-turn prompt templates based on the requirement and navigation route, filling them into predefined positions. Similarly, map images and the environment description are incorporated to further enrich the prompt content, providing strong support for VLM reasoning. These prompts are sequentially input into the VLM, enabling step-by-step reasoning through multi-turn conversations to produce intermediate outputs (e.g., scenario description) and final outputs (e.g., JSON format human behavior configuration). Each turn of conversation builds on previous context, ensuring the VLM continuously aligns with the test objective, understands the requirement, and generates the required output for each step.
After completing all conversation turns, RVSG delivers the final formatted output to the Configuration Decoder, which interprets the human configuration generated by the VLM based on the specific requirement and robot navigation route. The dialogue history is stored in the Memory module via Dialogue History Storage for future optimization. Moreover, the Scenario & Human Task Extractor retrieves the scenario description and human tasks (e.g., workers assigned to packing and delivery in a warehouse setting) from the dialogue, and summarizes this information to support experience accumulation for RVSG scenario generation.
The decoded human configuration is used to generate executable human agents in the simulator. To ensure realistic human behavior, RVSG employs HuNavSim [9], an agent controller based on behavior trees and the social force model, which closely simulates real-world human actions. The AMR under test is loaded with its designated navigation route and uses its navigation algorithm for environment perception and navigation decisions. Both human agents and the AMR perform their tasks in the simulator at 5, enabling dynamic performance evaluation of the AMR along the navigation route, with time series data collected throughout the process. The simulation environment is consistent with that used in preprocessing, and real-time communication is maintained via Robot Operating System 2 (ROS2) [10].
The time series data collected during simulation is used by the Objective Function Calculator to compute an objective function value, quantitatively reflecting the evaluation results based on the initial requirement. This result, together with the time series data, is submitted to the Feedback Integrator to generate feedback data. The feedback data can be used to optimize the original prompts in the Human Configuration Generation module for more effective scenario configuration generation. In addition, along with scenario and human task information, it is integrated into the Memory module through the Memory Integrator, which accumulates experience for VLM scenario generation, supporting prompt refinement and increased scenario diversity in future multi-turn conversations.
As shown at 6, with simulation feedback from VLM-generated test scenarios and scenario generation experience stored in the Memory module, the Prompt Generator in Human Configuration Generation can continually refine prompts. This enhances human configuration effectiveness for similar scenarios and increases the diversity of the generated test scenarios, enabling iterative testing. Ultimately, RVSG flexibly produces a more effective and diverse set of test scenarios based on simulation feedback, scenario diversity, and predefined stopping criteria (e.g., generating 10 scenarios), thereby improving the comprehensiveness and effectiveness of AMR testing.
3.2 Environment Preprocessing
RVSG uses a top-down view to capture the image of the simulation world, offering a clear representation of its structural and spatial relationships. To further highlight the environment layout, RVSG labels the image either manually or automatically by drawing bounding boxes around each area and assigning a corresponding name as a label. Since the VLM will analyze this labeled image and is highly adaptable in visual analysis, the labeling process is flexible and does not require strict standards, reducing the entry barrier for using RVSG. Based on the labeled map image, RVSG constructs a grid map with precise waypoints, with the adjustable resolution and size tailored to different maps and testing requirements. By default, RVSG uses evenly spaced square grids, where each grid point represents a unique 2D waypoint. For example, if rows are labeled with letters and columns with numbers, a specific waypoint, such as “E15”, refers to the point in row E and column 15. Only valid waypoints (i.e., those not obstructed) are included to aid route planning. In AMR testing, labeling is typically a one-time effort, allowing the same environment to be reused for testing different robots or for regression testing when their software is updated. As a result, manual labeling is more precise and not necessarily more costly than automated methods (e.g., AI methods), which could produce incorrect or imprecise results. Therefore, RVSG recommends using manual labeling methods for processing map and grid images.
Leveraging the VLM’s strong visual reasoning, RVSG conducts multi-turn conversations to analyze step-by-step the labeled map and grid images, extracting layout and navigation information. As shown in the Environment Description Generation module of Fig. 2, RVSG first provides the labeled map image and layout analysis instructions to the VLM to obtain a detailed analysis of environmental layout. Next, using the preserved dialogue history, RVSG sends the grid image and corresponding grid analysis instructions to the VLM, along with the map image that emphasizes the basic environmental information, and generates descriptions of the grid and navigation layouts. Finally, RVSG provides a summarization prompt to the VLM asking to summarize the key description for both map and grid images in JSON format. This output is then processed by the Environment Description Extractor, which extracts and pairs the corresponding descriptions for each image, storing them together as the current environment description for subsequent testing and analysis. Given the potential for hallucinations or inaccuracies in the VLM’s output, the final environment description is manually reviewed to ensure correctness.
3.3 Test Scenario Generation
As shown at 3 and 4 in Fig. 2, RVSG utilizes multi-turn conversations to generate scenarios and human configurations based on system requirements, robot navigation routes, and environment descriptions. By incorporating feedback from scenario simulations and memory of previous scenarios (at 6), RVSG further optimizes configuration generation and enhances scenario diversity. Specifically, RVSG provides multiple prompt templates for five key reasoning steps: (1) Test Scenario Description Generation, (2) High-level Human Task Generation, (3) Low-level Human Waypoint Path Generation, (4) Human Configuration Validity Check and Revision, and (5) JSON Format Human Configuration Generation.
Test Scenario Description Generation analyzes the current environment and produces realistic textual descriptions of test scenarios involving human activities. These descriptions serve as the basis for subsequent human configuration. High-level Human Task Generation allocates specific human tasks according to the generated scenario description, detailing concrete workflows and outputting the names of key labeled areas involved in these workflows. Low-level Human Waypoint Path Generation builds on the high-level tasks, sequentially selecting valid waypoints from the preprocessed grid map based on the workflow and key areas to construct human waypoint paths. The high-level human tasks and low-level waypoint paths together form the human configuration, which is then checked for validity by the Human Configuration Validity Check and Revision. Any problematic configurations are automatically detected and revised by the VLM at this stage. For example, the module verifies whether the workflow aligns with the scenario description and whether the waypoint paths are feasible. If inconsistencies are found, the workflow is modified and the waypoint paths are updated accordingly. Finally, JSON Format Human Configuration Generation provides the final human configuration in JSON format.
Fig. 3 presents all seven prompt templates provided by RVSG. Each template focuses on one or two specific key reasoning step(s). Notably, for Test Scenario Description Generation, the module has introduced an additional prompt version (Prompt2 in Fig. 3) beyond the initial prompt (Prompt1), influenced by previous scenario memory. Similarly, to regenerate the human configuration based on simulation feedback without modifying the underlying test scenario, RVSG integrates the prompts for High-level Human Task Generation and Low-level Human Waypoint Path Generation (Prompt3 and Prompt4), and provides a new Human Configuration Regeneration prompt (Prompt5). Based on these templates, RVSG automatically fills in the provided information, such as requirements, into the corresponding fields to generate prompts for VLM reasoning and analysis. For example, Prompt1 in Fig. 3 contains three fields: Environment Description, Robot Navigation Route, and Requirement. RVSG automatically populates each field with relevant content and outputs the final prompt. Note that the content for fields, such as Requirement and Robot Navigation Route, does not adhere to a strict format and can be flexibly adjusted based on experience.
Figure 3: Prompt templates for multi-turn conversations and scenario generation workflow with feedback and memory in RVSG.
RVSG combines different prompts to support iterative testing across various stages. As shown in Fig. 3, during the initial stage, RVSG selects the initial Test Scenario Description Generation prompt (Prompt1), High-level Human Task Generation prompt (Prompt3), Low-level Human Waypoint Path Generation prompt (Prompt4), Human Configuration Validity Check and Revision prompt (Prompt6), and JSON Format Human Configuration Generation prompt (Prompt7). These are combined to form the initial multi-turn conversation prompts that guide the VLM in generating the initial human configuration, which is then executed in the simulator to obtain simulation results. Subsequently, the feedback trigger determines whether to initiate feedback-based optimization. If optimization continues, RVSG combines the dialogue history with Human Configuration Regeneration (Prompt5), Human Configuration Validity Check and Revision (Prompt6), and JSON Format Human Configuration Generation (Prompt7) to guide the VLM in refining the human configuration based on simulation feedback, without altering the initial scenario description. The optimized configuration is then executed in the simulator, after which new feedback is obtained. This feedback includes the objective function value corresponding to the testing requirement, which is used to evaluate the effectiveness of the optimized configuration and determine the degree of optimization. If further optimization is needed, the generated feedback can be used iteratively in subsequent optimization steps. If the feedback trigger indicates stopping feedback optimization (e.g., reaching a certain number of optimizations), RVSG then relies on the memory trigger to decide whether to use memory for generating more diverse scenarios. If memory-based optimization proceeds, RVSG replaces the initial Test Scenario Description Generation prompt (Prompt1) with the Diversity Test Scenario Description Generation prompt (Prompt2), keeping the other prompts unchanged to guide the VLM in creating the new test scenario. Moreover, in Prompt2, the Historical Scenario Memory field provides the descriptions of all previously generated test scenarios, along with their corresponding optimal configurations and objective function values, serving as references for generating more diverse scenarios. Otherwise, the process terminates, and all generated scenarios are output. All multi-turn conversation histories, simulation results, and the corresponding scenario descriptions and human tasks are stored in the Memory module, supporting VLM-based scenario generation at different stages. Note that, throughout all multi-turn conversations, RVSG consistently provides labeled map and grid images either directly or indirectly. For simplicity, these details are omitted in the workflow shown in Fig. 3.
4 Experiment Design
4.1 Subject System and Simulator
We test the latest logistics AMR robot, TIAGo OMNI Base from PAL Robotics [11]. It has four mecanum wheels for omnidirectional movement and two LiDAR scanners that provide a 360° unobstructed field of view, enabling it to detect and autonomously avoid obstacles to ensure safe navigation. To support simulation and testing, PAL Robotics offers an open-source TIAGO OMNI Base ROS 2 Simulation [12]. The simulator is built on Gazebo classic 11 [13] and provides a high-fidelity model of the TIAGo OMNI Base. The TIAGo OMNI Base adopts Nav2 [14] for both global navigation and local obstacle avoidance, while real-time communication is achieved through ROS 2 Humble [10]. Considering the simulation platform and the ROS 2 version, we selected the compatible HuNavSim [9] as the default agent for human behavior control. At the group level, HuNavSim uses the social force model [15] to simulate general crowd flow and avoidance behaviors in the presence of obstacles or other humans. At the individual level, it employs behavior trees to define and control more complex, personalized human actions. As a result, HuNavSim can provide realistic human behavior modeling, which benefits AMR navigation task testing and supports the development of navigation systems.
4.2 Experiment Setup
S&F Requirements and Objective Functions. We consider three requirements provided by PAL Robotics: collision avoidance, stability, and efficiency. These requirements reflect essential safety and functional aspects of AMR navigation for PAL robotics and are listed as follows:
  • R1: Collision Avoidance – The robot must maintain a safe distance from other objects to prevent collisions.
  • R2: Stability – The robot must maintain operational stability, ensuring control is not lost due to external forces or loads.
  • R3: Efficiency – The robot should complete its assigned navigation task within a reasonable time, minimizing delays while maintaining safety and stability.
R1 is critical, as the risk of collision is a primary safety concern. To quantify the degree of violation of R1, we opt Distance to Obstacles (DTO[16] as the objective function:
$DTO=\frac{1}{T}\sum_{t=1}^{T}\min_{i=1}^{n}\left\|\mathbf{p}_{\text{robot}}^{t}-\mathbf{p}_{o_{i}}^{t}\right\|_{2},$ (1)
where $\mathbf{p}_{\text{robot}}^{t}$ and $\mathbf{p}_{o_{i}}^{t}$ are positions of the robot and object $o_{i}$ at timestep $t$, $n$ is the number of objects in the environment, and $T$ is the total number of timesteps. DTO measures the average minimum Euclidean distance from the robot to the nearest object during its navigation, where smaller values indicate a higher risk of collision.
R2 is also important, as a lack of stability may result in loss of control over the robot. To evaluate the violation of this requirement, we calculate Jerk [17] as the objective function:
$Jerk=\frac{1}{T-1}\sum_{i=1}^{T-1}\frac{\mathbf{a}_{i+1}-\mathbf{a}_{i}}{\Delta t},$ (2)
where $\mathbf{a}_{i}$ is the acceleration of the robot at timestep $i$, $\Delta t$ is the time interval between steps, and $T$ is the total number of timesteps. Jerk measures the rate of change of acceleration over time; higher jerk values indicate more abrupt motion changes, indicating unstable behaviors.
R3 concerns the time cost required to complete an operational task. To assess the violation of this requirement, we use Time-to-Reach-Goal (TRG) as the objective function:
$TRG=t_{\text{goal}}-t_{\text{start}},$ (3)
where $t_{start}$ and $t_{goal}$ are the timestamps when the robot starts moving and reaches its goal. TRG measures the total duration taken by the robot to finish a given navigation route. Longer TRG indicates inefficiencies or suboptimal behaviors in the robot’s execution of the task.
VLM and Parameter Settings. We employ GPT-4.1 [18] as the VLM and integrate it into RVSG for scenario generation. GPT-4.1 is the latest VLM from OpenAI’s GPT series, which is designed for complex tasks. It excels at performing a variety of tasks, including coding and instruction following. Moreover, it supports up to 1 million contextual tokens and can better leverage these contexts, improving long-context understanding ability. The parameter settings of VLMs can significantly influence the performance. In our setup, we adopt the default settings from OpenAI for GPT-4.1.
RVSG Feedback and Memory Settings. Based on a pilot study, we set the stopping condition for the feedback trigger to terminate after four configuration optimizations for the same scenario description, resulting in a total of five similar but distinct configurations per scenario. The stopping condition for the memory trigger was set to terminate after generating 10 sets of scenarios, each consisting of one scenario description and its corresponding optimal human configuration. Among these, the scenarios with the best and worst objective value were identified and repeated 30 times for further analysis.
Comparison Baselines. Arcuri and Briand suggested [19] to use random search (RS) to check if a problem is complex enough to be solved by a complex algorithm. Thus, we performed a pilot study to compare RVSG with RS. Results in our online repository [20] show that RVSG is significantly better than RS in generating scenarios that violate a given requirement. Hence, our formal experiment excluded RS as a baseline. Instead, we employ a modified RVSG as the comparison baseline, i.e., $\textit{RVSG}_{R}$. $\textit{RVSG}_{R}$ shares the same architectural design as RVSG but excludes the feedback and memory components, resulting in unguided scenario generation by the VLM.
Figure 4: Warehouse simulation world and navigation routes for AMR testing. Circles labeled with “S" and “E" indicate the starting and end of the routes, respectively.
Navigation Routes. We used the Amazon warehouse simulation world [21] as the test environment for our experiments. As shown in Fig. 4, the environment consists of multiple functional areas, with shelves, boxes, and other obstacles distributed across different regions, providing a complex and realistic simulation for AMR testing. We selected five different navigation routes that cover the entire warehouse. Some routes represent safer navigation scenarios. For example, in Route_1, the robot travels primarily through open areas with almost no nearby obstacles. Other routes include more challenging situations. For instance, in Route_3, the robot navigates very close to boxes and other obstacles in certain segments. This diverse set of routes enables a more comprehensive evaluation of the effectiveness of the generated test scenarios.
4.3 Evaluation Metrics
Objective Metric. We adopted the same three objective functions as in Section 4.2, i.e., DTO, Jerk, and TRG, as the objective evaluation metrics to directly assess the effectiveness of the generated test scenarios in violating different requirements.
Performance Metric. We incorporated four metrics from [9] to provide a performance assessment of the scenarios generated by RVSG for various requirement violations:
  • Robot on Person Collision (RPC): count the number of times the robot collides with a person. This metric is calculated using the positions, orientations, and velocities of both the robot and humans to determine collision events and identify which party initiates the collision.
  • Cumulative Heading Changes (CHC): measures the cumulative heading changes along the robot’s navigation, calculated as the sum of angles between successive waypoints. It reflects both path smoothness and motion stability, with higher values indicating less smooth and less stable movement.
  • Path Length (PL): measures the distance the robot travels from the initial point to the goal position, which is used to assess navigation efficiency, with longer paths generally indicating less efficient planning.
  • Time Not Moving (TNM): quantifies the seconds the robot remains stationary, which is used to evaluate the efficiency and continuity of the robot’s movement, with higher values indicating less continuous and effective navigation.
Diversity Metric. To evaluate the diversity of the scenarios generated by RVSG, we defined the following metrics:
  • Scenario Description Diversity (SDD): measures the diversity of natural language descriptions generated for scenarios, defined as one minus the average similarity computed by Sentence Transformers [22]. Higher values indicate greater diversity in scenario descriptions.
  • High-level Task Diversity (HTD): measures the diversity of human high-level task descriptions by computing the average Levenshtein distance [23] between them, which quantifies how many single-character edits (insertions, deletions, or substitutions) are needed to change one string into another. Higher values indicate greater diversity.
  • Scenario Simulation Diversity (SSD): measures the diversity of simulated scenarios by computing the average dynamic time warping (DTW) distance [24] between the time-series position data of the robot and humans. DTW calculates the distance between two temporal sequences, with larger distances indicating greater dissimilarity. Thus, higher values reflect greater diversity among simulated scenarios.
4.4 Research Questions
We addressed the following research questions (RQs): RQ1: How does RVSG perform compared to $\textit{RVSG}_{R}$ on different navigation routes? RQ2: How do RVSG-generated scenarios differ in exposing variability and instability in robot behavior? RQ3: How do navigation routes affect the performance and variability of RVSG-generated scenarios?
4.5 Statistical Tests
Since our data are continuous and may not satisfy the normality assumption, we chose the following guide [25] to select the Mann-Whitney U test for determining statistical significance of our results (significance level 0.05). Effect sizes were assessed with the Vargha and Delaney Â12 statistic, with values of $\geq$ 0.638 or $\leq$ 0.362 considered indicative of medium or large effect sizes based on the guides [25, 26].
5 Results and Analysis
5.1 Comparison between RVSG and $\textit{RVSG}_{R}$— RQ1
Table 1: Results achieved by RVSG and $\textit{RVSG}_{R}$ for all metrics on different routes and requirements. “A/B" means the results from $\textit{RVSG}_{R}$ and RVSG. A is from $\textit{RVSG}_{R}$ and B is from RVSG. Bold indicates a method with better performance. – RQ1
Requirement
Route Objective Metric Performance Metric Diversity Metric
RPC $\uparrow$ CHC $\uparrow$ PL $\uparrow$ TNM $\uparrow$ SDD $\uparrow$ HTD $\uparrow$ SSD $\uparrow$
Collision
Avoidance
(DTO $\downarrow$)
Route_1 3.020 / 0.666 0.2 / 3.7 2.002 / 2.222 8.482 / 8.491 0.249 / 0.299 0.142 / 0.253 4.550 / 4.950 60.415 / 41.385
Route_2 3.717 / 0.809 0.1 / 1.9 1.345 / 2.072 9.279 / 9.398 0.299 / 0.351 0.127 / 0.209 4.744 / 6.244 70.531 / 39.966
Route_3 1.950 / 0.390 0.7 / 8.0 2.275 / 2.837 8.136 / 8.934 1.849 / 0.250 0.136 / 0.234 4.611 / 4.783 55.667 / 27.019
Route_4 3.433 / 1.432 1.2 / 1.3 1.036 / 2.034 10.996 / 11.118 0.301 / 0.351 0.148 / 0.187 4.067 / 6.228 56.523 / 51.388
Route_5 3.226 / 1.501 3.5 / 5.3 1.386 / 1.969 14.411 / 14.656 0.201 / 0.200 0.127 / 0.228 4.217 / 5.928 69.550 / 51.151
Stability
(Jerk $\uparrow$)
Route_1 0.980 / 1.371 0.2 / 0.3 2.002 / 2.020 8.482 / 8.472 0.249 / 0.248 0.142 / 0.200 4.550 / 8.228 60.415 / 61.838
Route_2 0.613 / 1.354 0.1 / 1.0 1.345 / 1.813 9.279 / 9.366 0.299 / 0.449 0.127 / 0.200 4.744 / 8.522 70.531 / 61.546
Route_3 0.949 / 2.239 0.7 / 1.8 2.275 / 3.051 8.136 / 8.941 1.849 / 0.399 0.136 / 0.188 4.611 / 8.706 55.667 / 45.093
Route_4 0.890 / 2.698 1.2 / 0.9 1.036 / 1.427 10.996 / 11.049 0.301 / 0.150 0.148 / 0.181 4.067 / 9.133 56.523 / 60.108
Route_5 1.636 / 1.221 3.5 / 1.0 1.386 / 1.311 14.411 / 14.441 0.201 / 0.200 0.127 / 0.200 4.217 / 7.956 69.550 / 71.218
Efficiency
(TRG $\uparrow$)
Route_1 20.433 / 21.024 0.2 / 0.4 2.002 / 2.249 8.482 / 8.477 0.249 / 0.350 0.142 / 0.183 4.550 / 9.461 60.415 / 53.401
Route_2 21.585 / 23.024 0.1 / 0.2 1.345 / 2.444 9.279 / 9.507 0.299 / 0.249 0.127 / 0.222 4.744 / 11.417 70.531 / 71.965
Route_3 21.695 / 22.066 0.7 / 0.6 2.275 / 2.615 8.136 / 8.822 1.849 / 0.299 0.136 / 0.186 4.611 / 11.683 55.667 / 60.087
Route_4 23.958 / 24.606 1.2 / 0.8 1.036 / 1.506 10.996 / 11.016 0.301 / 0.150 0.148 / 0.190 4.067 / 7.717 56.523 / 64.172
Route_5 31.189 / 33.217 3.5 / 5.8 1.386 / 2.194 14.411 / 14.505 0.201 / 0.300 0.127 / 0.180 4.217 / 9.956 69.550 / 78.784
Table 1 shows the results of RVSG and $\textit{RVSG}_{R}$ across all metrics and navigation routes. We denote RVSG under three requirement settings as $\textit{RVSG}_{\textit{DTO}}$, $\textit{RVSG}_{\textit{Jerk}}$, and $\textit{RVSG}_{\textit{TRG}}$. When looking at Objective Metric, for R1, $\textit{RVSG}_{\textit{DTO}}$ consistently outperformed $\textit{RVSG}_{R}$ across all routes in DTO, indicating more effective violation of R1 by RVSG than $\textit{RVSG}_{R}$. For R2, $\textit{RVSG}_{\textit{Jerk}}$ led to higher jerk on four routes (i.e., Route_1-4), showing better violation of R2 by RVSG than $\textit{RVSG}_{R}$. For R3, $\textit{RVSG}_{\textit{TRG}}$ outperformed $\textit{RVSG}_{R}$ on all routes, meaning $\textit{RVSG}_{\textit{TRG}}$ generated scenarios that effectively hindered the robot from reaching its goal.
For Performance Metric, $\textit{RVSG}_{\textit{DTO}}$ outperformed $\textit{RVSG}_{R}$ in RPC, showing that violating R1 increases the chances of robot-responsible collisions. However, $\textit{RVSG}_{\textit{Jerk}}$ and $\textit{RVSG}_{\textit{TRG}}$, led to higher RPC in only about half of the cases, since R2 and R3 are not related to collision and thus provide weaker guidance for collision than $\textit{RVSG}_{\textit{DTO}}$. For CHC and PL, RVSG outperformed $\textit{RVSG}_{R}$ in 27 out of 30 cases across all requirements and routes. In the remaining three cases, differences were marginal (i.e., 1.386 vs. 1.311, 8.482 vs. 8.472, 8.482 vs. 8.477), indicating that RVSG consistently generates scenarios that effectively degrade navigation performance. For TNM, the performance between RVSG and $\textit{RVSG}_{R}$ was comparable; in certain cases, RVSG performed better, while in others, it performed slightly worse. Scenario replay showed that higher TNM does not always reflect worse performance, as it can also represent appropriate pauses or safe waiting behaviors.
For Diversity Metric, for SDD and HTD, RVSG outperformed $\textit{RVSG}_{R}$ across all requirements and routes, showing greater diversity in scenario descriptions and high-level task descriptions. However, for SSD, RVSG outperformed $\textit{RVSG}_{R}$ in 7 of 15 cases (0 for R1, 3 for R2, 4 for R3). This suggests that RVSG’s strength lies in linguistic and task-level diversity rather than in producing diverse human-robot interactions. A possible reason is that $\textit{RVSG}_{R}$, resembling random search, explores the simulation space more broadly.
Conclusion for RQ1: RVSG outperformed $\textit{RVSG}_{R}$ in generating scenarios that more effectively violate safety and performance requirements, degrading robot performance, and enhancing linguistic and task-level diversity. However, RVSG’s ability to simulate diverse scenarios is limited, likely due to its more guided search strategy.
5.2 Assessing Robot Behavioral Variability and Instability with RVSG-Generated Scenarios— RQ2
Table 2: Statistical test, standard deviation, and diversity results for the best and worst scenarios generated by RVSG (repeated 30 times) across different metrics and requirements. “/" denotes no statistical results. Bold indicates a significant result. – RQ2
Metric Collision Avoidance (R1) Stability (R2) Efficiency (R3)
p Â12 $\textit{RVSG}_{B}$ $\textit{RVSG}_{W}$ p Â12 $\textit{RVSG}_{B}$ $\textit{RVSG}_{W}$ p Â12 $\textit{RVSG}_{B}$ $\textit{RVSG}_{W}$
DTO $\bm{<}$0.01 0.238 0.776 0.656 $\bm{<}$0.01 0.253 0.496 0.494 $\bm{<}$0.01 0.242 0.315 0.375
Jerk $\geq$0.05 0.507 1.834 1.445 $<$0.01 0.615 2.296 0.902 $<$0.01 0.602 0.584 0.809
TRG $\geq$0.05 0.504 1.014 0.857 $<$0.05 0.581 0.943 0.381 $<$0.01 0.595 1.041 0.594
RPC $<$0.05 0.579 5.579 2.734 $<$0.01 0.582 1.150 0.392 $<$0.01 0.599 1.777 1.611
CHC $<$0.05 0.568 0.921 0.693 $\bm{<}$0.01 0.710 0.625 0.232 $\bm{<}$0.01 0.797 0.737 0.294
PL $\geq$0.05 0.547 0.160 0.097 $\geq$0.05 0.546 0.323 0.065 $\geq$0.05 0.560 0.151 0.053
TNM $\geq$0.05 0.555 0.277 0.298 $\geq$0.05 0.477 0.712 0.264 $\geq$0.05 0.520 0.266 0.252
SSD / / 22.067 21.556 / / 17.690 18.973 / / 21.559 13.138
We picked the best and worst scenarios from all the scenarios generated by RVSG. We executed each selected scenario 30 times to study its characteristics for variability in the robot’s behaviors. For Objective Metric and Performance Metric, we performed the Mann-Whitney U test and the Vargha and Delaney Â12 effect size to compare differences between all the best ($\textit{RVSG}_{B}$) and worst ($\textit{RVSG}_{W}$) scenarios. We also calculated the standard deviation (std) for these metrics to understand the variability in the robot’s behavior across repeated executions. For Diversity Metric, we calculated SSD across the 30 simulations for each scenario. Table 2 summarizes the results for different requirements. For R1, results showed no statistically significant difference between $\textit{RVSG}_{B}$ and $\textit{RVSG}_{W}$ in all metrics except for DTO, indicating similar robot performance. Nevertheless, for DTO, $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$, meaning it better violated R1. Moreover, $\textit{RVSG}_{B}$ consistently achieved higher std values for almost all metrics. This indicates less stable robot behaviors when using $\textit{RVSG}_{B}$ compared to $\textit{RVSG}_{W}$, suggesting that $\textit{RVSG}_{B}$ is more suitable for uncovering unexpected behaviors during testing. For R2, $\textit{RVSG}_{B}$ achieved a comparable performance as $\textit{RVSG}_{W}$ for five out of seven metrics (i.e., Jerk, TRG, RPC, PL, TNM), and for the other two metrics (DTO, CHC), $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$. Regarding variability, $\textit{RVSG}_{B}$ consistently achieved higher std values across all metrics, indicating less stable robot behaviors during testing. Similarly, for R3, $\textit{RVSG}_{B}$ achieved comparable or significantly better performance than $\textit{RVSG}_{W}$ across the metrics, while also leading the robot to exhibit more unstable behaviors. This consistent pattern indicates that the best scenarios identified by RVSG are more effective at exposing unexpected or unstable behaviors of the robot, making them more suitable for testing. Finally, the diversity results (i.e., SSD) show that $\textit{RVSG}_{B}$ performed more unstable than $\textit{RVSG}_{W}$ for R1 and R3, while for R2, $\textit{RVSG}_{B}$ is more stable than $\textit{RVSG}_{W}$. This variation may lie in the requirements, and further experiments are needed to confirm this hypothesis.
Conclusion for RQ2: $\textit{RVSG}_{B}$ induces more unstable robot behaviors than $\textit{RVSG}_{W}$ across different requirements. Besides, $\textit{RVSG}_{B}$ is consistently more effective than $\textit{RVSG}_{W}$ and, in several cases, significantly outperforms $\textit{RVSG}_{W}$ on certain metrics such as DTO and CHC.
5.3 Studying the Effect of Navigation Routes in RVSG-Generated Scenarios— RQ3
We further investigated the impact of different routes on the performance of the test scenarios generated by RVSG with same settings as RQ2 (i.e., picking best and worst scenario and the same statistics). We summarize the key findings here, and all detailed results are available in our public repository [20].
For R1, $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$ for DTO on four out of the five routes, with the only exception being Route_5. On Route_5, no significant differences were observed across all metrics. Notably, Route_5 is a straight path without any turns, resulting in a lower task complexity. This may diminish the performance differences between the two scenarios, leading to the absence of significant differences in the results. For other metrics, $\textit{RVSG}_{B}$ was only significantly worse than $\textit{RVSG}_{W}$ in a few cases. However, changes in the navigation route could have also affected the results. For example, in Route_1, $\textit{RVSG}_{B}$ performed significantly worse than $\textit{RVSG}_{W}$ for Jerk, while in Route_4, it performed better. This may be because the generated test scenarios, combined with routes passing close to shelves and boxes (as in Route_4), introduce additional complexities. These factors further restrict the robot’s movement space and require more frequent adjustments, thereby impacting the Jerk values. In addition, we observe that in Route_3 and Route_4, $\textit{RVSG}_{B}$ had a higher std than $\textit{RVSG}_{W}$. For other routes, the std for different metrics varied. This suggests that the navigation route affects the stability of scenario execution and leads to different levels of variability. For R2, $\textit{RVSG}_{B}$ performed significantly worse than $\textit{RVSG}_{W}$ in only two metrics (DTO and CHC) in Route_1. However, in other routes, $\textit{RVSG}_{B}$ significantly outperformed $\textit{RVSG}_{W}$ in terms of these two metrics. Regarding the std, $\textit{RVSG}_{B}$ showed higher values than $\textit{RVSG}_{W}$ in some routes and lower values in others. This suggests that the navigation route influences the performance of scenarios generated by RVSG. For R3, $\textit{RVSG}_{B}$ outperformed $\textit{RVSG}_{W}$ or showed no significant difference in most metrics. The std of $\textit{RVSG}_{B}$ was higher than $\textit{RVSG}_{W}$ in some routes and lower in others, indicating that the route affects the performance of RVSG. For SSD, $\textit{RVSG}_{B}$ exhibited greater diversity than $\textit{RVSG}_{W}$ in some routes but less in others, suggesting influence by the route. Fig. 5 shows the distribution results of $\textit{RVSG}_{B}$ and $\textit{RVSG}_{W}$ for different metrics and requirements across various routes. Overall, the navigation routes influenced the distribution of both scenarios for the same requirement or metric. Moreover, the figure indicates that in Route_3 and Route_4, $\textit{RVSG}_{B}$ exhibited higher distribution than $\textit{RVSG}_{W}$, suggesting that the robot displays greater behavioral variability in scenarios produced by $\textit{RVSG}_{B}$.
Figure 5: Distribution results of the best and worst scenarios (repeated 30 times) generated by RVSG with different requirements for different metrics under different navigation routes. The mean represents the central tendency. – RQ3
Conclusion for RQ3: Navigation routes affect the performance of $\textit{RVSG}_{B}$ and $\textit{RVSG}_{W}$ under different requirements and metrics, as well as the stability of scenario execution and the robot’s behavioral diversity. Different routes may lead to significant changes in metrics’ values and distributions, highlighting the crucial role of navigation routes.
6 Threats to Validity
Concerning conclusion validity concerns, to draw a reliable conclusion, we employed appropriate statistical tests and a rigorous analysis based on established guides. Regarding internal validity concerns related to parameters, we used the default parameters for the VLM from OpenAI and kept the same parameter settings for all approaches. Concerning constructive validity threat, we used human-aware navigation metrics from HuNavSim [9]. These metrics are comparable and commonly applied in robot navigation [27, 28]. To mitigate threats due to hallucinations and errors in VLM outputs, we manually reviewed a subset of the generated responses. External Validity concerns the generalizability of RVSG and the empirical evaluation. We employed three requirements and five robot navigation routes. The selected requirements represent key safety and functionality of the robot and are provided by PAL Robotics. The navigation routes are selected in consultation with PAL Robotics to reflect real-world operations and human-robot interactions. To address the simulator fidelity threat, we used Gazebo, a high-fidelity simulator that enables realistic modeling of robot dynamics, sensor noise, and human-agent interactions. Moreover, this simulator is used by PAL Robotics in their practice. Besides, to enhance the realism of human behaviors, we used HuNavSim [9], which integrates the social force model and diverse human reactions to robot presence.
7 Lessons Learned and Industrial Perspectives
Benefits of VLMs in Test Scenario Generation. VLMs offer several key benefits for automated scenario generation for AMR navigation testing. First, they can process multimodal inputs (e.g., visual and textual data) enabling them to generate environmental context-dependent scenarios (including human behaviors) that are logical for a given environment. For example, in a warehouse setting, a VLM can generate realistic human behaviors, such as workers picking items from shelves. Second, VLMs possess commonsense reasoning capabilities [6], allowing them to infer plausible risky human behaviors that could violate requirements. Third, their natural language understanding enables interpretation of high-level F&S requirements, ensuring that generated scenarios are focused on violating those requirements. Fourth, VLMs support diverse scenario generation, which enhances test coverage. With feedback and memory mechanisms, RVSG iteratively refines scenario quality, allowing learning from historical failures, adjusting human behavior configurations, and avoiding redundancy.
Practical Implications. The practical adoption of RVSG can facilitate the development and testing of industrial AMRs. RVSG enables cost-effective and scalable testing by automating the generation of realistic, requirement-targeted scenarios without significant human effort. For example, a tester can specify requirements in natural language, which can then directly guide test generation, reducing both the need for technical expertise in robot testing and the overall implementation effort for coding tests. VLMs are demonstrated for one testing application in RVSG; however, they can also support other testing activities, such as automated test oracle generation and regression testing. By interpreting high-level requirements and analyzing multimodal outputs, VLMs can help determine expected outcomes, thereby reducing manual effort to devise test oracles. Furthermore, VLMs can support AMR regression testing by automatically identifying updates in software across evolution and using them to guide regression testing (e.g., test case selection). Lastly, VLMs can support other AMR development phases, including requirements engineering (e.g., requirements identification), modeling and analysis (e.g., tools for automated creation of robot simulation models), and software development (e.g., automating code generation).
8 Related Works
Testing of Robotic Systems. Various robotic testing approaches have been proposed, e.g., simulation-based testing [29, 30, 31, 32] and formal verification [33, 34]. For example, Hutchison et al. [35] proposed a framework to automate the test generation and execution for industrial autonomy systems, including robots, and robotics-oriented libraries. Brito et al. [36] proposed an integration testing approach for robotic systems focusing on the publish/subscribe communication function and generating critical test scenarios guided by coverage criteria. To ensure the correct ROS operation, Santos et al. [32] presented a property-based testing approach for ROS system configurations, focusing on orchestration errors that occur when integrating different ROS components at runtime. Compared to the above methods, RVSG focuses on system-level testing of an industrial robot to generate scenarios including human behaviors in simulation. It uses a VLM to search for semantically meaningful and realistic scenarios that challenge the robot’s behavior to violate its functional and safety requirements.
Foundation Models-based Test Generation. Foundation models (FMs), e.g., LLMs and VLMs, have shown benefits in many areas. Among them, FM-based test generation offers an effective way to generating tests [37, 38, 39]. With techniques, e.g., prompting or fine-tuning, FMs support testing activities, including unit testing [39, 40] and test oracle generation [41, 42]. Though FMs showed good performance in testing tasks like unit testing, their use to test complex software systems, e.g., robots, remains unexplored due to their complexity (e.g., real-time constraints and multimodal inputs), which pose unique testing challenges. Some works address these challenges for autonomous vehicles with FMs [43, 44]. In contrast, RVSG generates test scenarios for an industrial robot operating in dynamic environments. Specifically, RVSG focuses on scenarios involving human-robot interactions, which are critical to both functionality and safety. Using VLMs, RVSG can synthesize realistic scenarios that expose potential violations of robots’ safety and functional requirements.
9 Conclusions and Future Work
We present RVSG, a VLM-based test scenario generation approach for robot navigation. Given a functional or safety requirement, RVSG generates scenarios violating the requirement. RVSG employs a VLM to extract visual features from an image describing a robot’s environment. Next, the extracted features are then used by the VLM to generate human behavior configurations along with the scenario description by iteratively querying the VLM with specifically designed prompts. The scenarios generated in each iteration are then executed in simulation, and the execution results are used as feedback for the next iteration to refine the scenarios. We evaluated RVSG in a warehouse environment and selected three requirements and five robot navigation routes for the experiment. The results show that RVSG can generate effective requirement-violating test scenarios that increase the variability of robot behaviors, helping reveal behavioral uncertainties. Meanwhile, the navigation routes affect the performance and stability of RVSG-generated scenarios. Future works include studying the generalizability of RVSG by conducting experiments with other robots. Also, we plan to investigate multi-objective RVSG that can violate multiple requirements simultaneously. Our open-source repository is now publicly available online [20].
Acknowledgments
This work is supported by the Co-tester project (No. 314544) and the Co-evolver project (No. 286898/F20), funded by the Research Council of Norway. The work is also supported by the RoboSAPIENS project funded by the European Commission’s Horizon Europe programme under grant agreement number 101133807. Aitor Arrieta is part of the Software and Systems Engineering research group of Mondragon Unibertsitatea (IT1519-22), supported by the Department of Education, Universities and Research of the Basque Country.